Alexis is selling tickets to the school play.Adult tickets cost $6.00,and student tickets cost $4.00. she sells twice as many students tickets as adult tickets.The total amount collected is $462.how many of each type of ticket did she sell?

Answer by checkley79(3341)

X IS THE NUMBER OF ADULT TICKETS SOLD.
2X IS THE NUMBER OF STUDENR TICKETS SOLD.
6X+4*2X=462
6X+8X=462
14X=461
X=461/14
X=33 ADULT TICKETS WERE SOLD.
2*33=66 STUDENT TICKETS WERE SOLD.
PROOF:
6*33+4*66=462
198+264=462
462=462